Project

General

Profile

NDOS Commissioning Meeting 2011/02/10

Attendance: Jon, Chad, Andrew, Peter, Gavin, Tricia, Mark Mathis, Minerba, Alena, Athans, Mat, Susan

Agenda:
  1. DAQ/Electronics Update & Plans
  2. DCS Status & Plans
  3. Shifters' report
  4. AOB

DAQ:

  • Peter:
    • We've been struggling with the 3 hour spike problem.
    • Andrew:
      • first came up on Jan. 27 when a change was made to TDU interface
      • did not realize that a change was also made to firmware
      • interface remapped a bit
      • resulted in autosyncs happening very frequently
      • it is thought that there's a bug in the autosync feature that
    • Peter is having a hard time starting a run; not so much with corrupt uslices, but uslices are out-of-time
    • Andrew: are time markers being enabled? If not, we may have to manually do it for now...
    • Other things coming down the pipeline:
      • configuration manager <-> dB
      • rollout of new version of DDS
      • status readback work has been put on hold due to lack of stability
    • Rick came up with version of DCM firmware to log timing commands
  • Mat (channel recovery):
    • 11 new APDs were installed on the first block
    • DSO scan was just done
    • blue tint was found on fibers of one module. Believed to have been cleaned.
  • DCM Spare Repair (Andrew):
    • dcm7 seems to be working, but it suddenly started working, so there's come concern that this is not understood.
    • Rick K would like another day or so to investigate.
    • will make a DCM available for NDOS installation tomorrow
  • DCM installation:
    • Rick T: need 3 DCMs in the next week
    • Move from DCM test-stand to NDOS tomorrow
    • need to develop DCM checkout procedure - Jon will ask Sue and Ron for advice
    • next week, will take one extra DCM off for installation at NDOS
    • another DCM will be taken from the test-stand and the 3rd di-block will use only 2 DCMs (will require non-trivial channel-map)
    • would be good to have checkout procedure for DCM->FEB communication
  • Channel map:
    • two channels are swapped
    • Leon will swap them back into proper order today
    • Jon will work on channel map exceptions dB table
  • Leon:
    • Hamamatsu is looking into using a thin-enough coating for our needs
    • Work is progressing on reducing TECC noise.
      • Various ways (copper tape, shielding in different places) have been found to kill the noise.
      • Will need to continue work tomorrow.
  • Rick (FEB installation)
    • muon catcher + secondary containment installation comes first (about 5-7 business days)
    • FEB installation will follow

DCS:

  • Chad:
    • a lot of temperatures alarms have been going off recently
    • initial theory was that heater in stairwell was culprit; turned space heater down, but racks are even hotter
    • thermostat is located at ground level.
    • Rick T. will look into getting fans installed to help circulate
    • also added an alarm history tab
  • Peter: dummy chillers are falling apart; will remove these tomorrow
  • Athans: committed first version of sender-receiver applications for DCM DCS, but code is not ready for use and probably won't be until March.

Any Other Business:

  • Mark Mathis: doc-db 5269, Analysis of DSO Data
    • reproduced what has been done before
    • found several disconnected APDs, but has only looked at older data (from a few weeks ago), so some of these channels may have been recovered.
    • Leon: it would be good to look at those APDs that go in and out of scans.
    • Andrew is working on making longer DSO scans...